# CamSAM2: Segment Anything Accurately in Camouflaged Videos

### v2

@@ Abstract
Video camouflaged object segmentation (VCOS), aiming at segmenting camouflaged objects that seamlessly blend into their environment, is a fundamental vision task with various real-world applications. With the release of SAM2, video segmentation has witnessed significant progress. However, SAM2’s capability of segmenting camouflaged videos is suboptimal, especially when given simple prompts such as point and box. To address the problem, we propose Camouflaged SAM2 (CamSAM2), which enhances SAM2’s ability to handle camouflaged scenes without modifying SAM2’s parameters. Specifically, we introduce a decamouflaged token to provide the flexibility of feature adjustment for VCOS. To make full use of fine-grained and high-resolution features from the current frame and previous frames, we propose implicit object-aware fusion (IOF) and explicit object-aware fusion (EOF) modules, respectively. Object prototype generation (OPG) is introduced to abstract and memorize object prototypes with informative details using high-quality features from previous frames. Extensive experiments are conducted to validate the effectiveness of our approach. While CamSAM2 only adds negligible learnable parameters to SAM2, it substantially outperforms SAM2 on three VCOS datasets, especially achieving 12.2 mDice gains with click prompt on MoCA-Mask and 19.6 mDice gains with mask prompt on SUN-SEG-Hard, with Hiera-T as the backbone. The code will be available at github.com/zhoustan/CamSAM2.
@@ 1 Introduction
Figure 1: Illustration of SAM2 and CamSAM2. (d) SAM2’s segmentation of the camouflaged object is suboptimal, primarily because its feature optimization is biased toward natural videos, and its design does not account for the unique challenges inherent to VCOS. (e) CamSAM2 improves SAM2’s ability on segmenting and tracking camouflaged objects by introducing a decamouflaged token, IOF to enhance features with high-resolution features, and EOF and OPG to further enhance features by exploiting informative object details across time. CamSAM2 only adds a limited number of parameters on SAM2 while keeping all SAM2’s parameters fixed and fully inheriting SAM2’s zero-shot ability.
@@
Camouflaged object detection (COD) and video camouflaged object segmentation (VCOS) aim to identify objects that blend seamlessly into their surroundings. Unlike standard object segmentation tasks, where objects typically exhibit clear boundaries and contrast with the background, camouflaged objects are naturally indistinguishable from the background. These tasks have various applications in wildlife monitoring, surveillance, and search-and-rescue operations [40, 44]. COD focuses on detecting camouflaged objects in individual images, while VCOS extends it to video sequences, adding the complexity of modeling temporal information across frames. Despite recent advancements in COD [33, 7, 58, 5, 34, 59, 54] and VCOS [33, 7, 15, 29, 20, 4, 25], the performance remains far from satisfactory compared to standard segmentation tasks.
@@
The recently introduced vision foundation model, Segment Anything Model 2 (SAM2) [39], marks a significant advancement in video segmentation. SAM2 has learned rich and generalizable representations for natural scenes on the SA-1B [19] (11M images, 1B masks) and SA-V [39] (50.9K videos, 35.5M masks) datasets. Therefore, its features are optimized for natural scenes, while SAM2’s ability of segmenting camouflaged objects is suboptimal, as in [61, 47]. As shown in Fig. 1, SAM2 segments only part of a camouflaged animal (hedgehog) with a single-click prompt, indicating that there is still room for performance improvements in VCOS using SAM2.
@@
This paper aims to develop a model for accurate segmentation in camouflaged videos, requiring both natural image understanding and effective identification of camouflaged objects in complex environments. To achieve this, we identify the following core challenges in adapting SAM2 for VCOS: (1) SAM2 is optimized for natural scenes rather than camouflaged environments. (2) The architecture does not account for the complexities of segmenting and tracking camouflaged objects across time. For VCOS, accurately segmenting camouflaged objects for a frame requires: a) exploiting fine-grained and detailed features from the frame, and b) considering the temporal evolvement of fine-grained features from previous frames. For exploiting temporal information, SAM2 is equipped with a memory module containing a memory encoder and memory bank. However, only low-resolution and coarse features are encoded into the memory, which is suboptimal for accurate VCOS.
@@
To tackle the above limitations and fully keep SAM2’s ability to process natural videos, we introduce Camouflaged SAM2, dubbed as CamSAM2, equipping SAM2 with the ability to effectively tackle VCOS, as depicted in Fig. 1. CamSAM2 includes a learnable decamouflaged token, which provides flexibility to optimize features for VCOS without modifying SAM2’s trained parameters. To exploit the fine-grained features of the frame, we propose the Implicit Object-aware Fusion (IOF) module, which enhances features with implicitly object-aware information. To make use of detailed features from previous frames, we further propose Object Prototype Generation (OPG) to abstract high-quality features within the object region into informative object prototypes through Farthest Point Sampling (FPS) and k-means. Those object prototypes are saved to the memory for easy usage by the Explicit Object-aware Fusion (EOF) module that is designed to integrate explicit object-aware information across the temporal dimension. Our design avoids saving the high-resolution features in the memory and only adds negligible computations to SAM2 while accounting for a large amount of temporal information.
@@
We conduct extensive experiments in §4 to validate the effectiveness of CamSAM2 on three VCOS benchmarks: two camouflaged animal datasets, MoCA-Mask [7] and CAD [3]; and one camouflaged medical dataset, SUN-SEG [17]. Our experiments show that CamSAM2 significantly outperforms SAM2 by achieving improvements of 12.2/13.0 mDice scores with click prompt on MoCA-Mask for Hiera-T/Hiera-S backbones, and 19.6 mDice gains with mask prompt on SUN-SEG-Hard for Hiera-T backbone. When directly evaluating CamSAM2 on CAD without further finetuning, we observe strong zero-shot ability. Since all SAM2’s weights remain unchanged, CamSAM2 totally inherits SAM2’s capability on segmenting natural videos. In summary, our contributions are three-fold:
  • We propose CamSAM2 to equip SAM2 with the ability to segment and track camouflaged objects in videos while keeping SAM2’s strong generalizability in natural videos.
@@
  • CamSAM2 introduces a decamouflaged token to achieve easy feature adjustments for the VCOS task without affecting SAM2’s trained weights. To effectively exploit the crucial fine-grained and high-resolution features from both the current frame and previous frames, we propose IOF, EOF, and OPG modules.
  • Our approach clearly outperforms SAM2 and sets new state-of-the-art performance on public VCOS datasets. Experiments also show the strong zero-shot ability of CamSAM2 in the domain of VCOS.
@@ 2.1 Camouflaged Object Detection
Camouflaged Scene Understanding (CSU) focuses on interpreting scenes where objects blend closely with their backgrounds, such as natural environments like forests, oceans, and deserts. Early works in this field primarily involved the collection of extensive image and video datasets, such as CAMO [23], COD10K [10], NC4K [26], CAD [3], MoCA-Mask [7], and MoCA-Mask-Pseudo [7], which laid the foundation for CSU.
Traditional Camouflaged Object Detection (COD) methods extract foreground-background features using optical features [2], color, and texture [18]. Deep learning has advanced COD with CNNs and transformers. SINet [10] and SINet-V2 [12] enhance fine-grained cues by applying receptive fields and texture-enhanced modules, while DQNet [45] applies cross-modal detail querying to detect subtle features. Transformer-based models like CamoFormer [54] leverage multi-scale feature extraction with masked separable attention, and WSSCOD [58] employs a frequency transformer and noisy pseudo labels for weak supervision. ZoomNeXt [33] further optimizes multi-scale extraction via a collaborative pyramid network. These advancements refine COD by integrating sophisticated architectures and diverse supervision strategies.
@@ 2.2 Video Camouflaged Object Segmentation
VCOS [25, 57, 20, 49] extends COD to videos, introducing challenges from motion, dynamic backgrounds, and temporal consistency. Former VCOS models tackle these with motion learning, spatial-temporal attention, and advanced segmentation techniques to maintain object coherence across frames. Motion-guided models enhance segmentation by leveraging motion cues. IMEX [14] integrates implicit and explicit motion learning for robust detection. TMNet [57] refines motion features with a transformer-based encoder and neighbor connection decoder. Flow-SAM [50] uses optical flow as input or a prompt, guiding SAM to detect moving camouflaged objects. Spatial-temporal attention enhances the tracking of camouflaged objects. TSP-SAM [15] and SAM-PM [29] improve SAM’s ability to detect subtle movements. Static-Dynamic-Interpretability [20] quantifies static and dynamic information in spatial-temporal models, aiding balanced approaches. Assessing camouflage quality is also essential for VCOS. CAMEVAL [22] introduces scores evaluating background similarity and boundary visibility, refining datasets and improving model robustness. These advancements drive more accurate and effective VCOS systems.
@@ 2.3 Segment Anything Model 2
SAM2 [39] is a vision foundation model for promotable segmentation across images and videos. Compared to SAM [19] which is limited to image segmentation, SAM2 offers a significant performance leap in video segmentation, producing higher segmentation accuracy while using fewer interactions than previous methods. SAM2 has demonstrated strong capabilities in many tasks, including medical image, video and 3D segmentation [27, 42, 56, 24, 13, 6, 51, 55], video object tracking and segmentation [43, 32], remote sensing [38], 3D mesh and point cloud segmentation [46], image camouflaged object detection and video camouflaged object segmentation [61, 6, 51, 47]. In previous works [61, 47], although SAM2’s ability to segment camouflaged videos has surpassed most existing methods, there is still a significant performance gap compared to other VOS tasks, especially when using simple prompts.
@@ 3 Method
Figure 2: Overall architecture of CamSAM2. CamSAM2 effectively captures and segments camouflaged objects by leveraging implicit and explicit object-aware information from the current or previous frames. It includes the following key components: (a) the decamouflaged token, which extends SAM2’s token structure to learn features suitable for segmenting camouflaged objects; (b) an IOF module to enrich memory-conditioned features with implicitly object-aware high-resolution features; (c) an EOF module to aggregate explicit object-aware features; and (d) an OPG module, generating informative object prototypes, which guides cross-attention in EOF. These components work together to preserve fine details, enhance segmentation quality, and track camouflaged objects across time.
We propose CamSAM2, equipping SAM2 with the ability to accurately segment camouflaged objects in videos while retaining SAM2’s original capabilities. §3.1 briefly reviews the architecture of SAM2. From §3.2 to §3.5, we describe CamSAM2 tailored for VCOS. With fixing SAM2’s parameters, CamSAM2 proposes a learnable decamouflaged token, Implicit and Explicit Object-aware Fusion, and Object Prototype Generation to enhance feature representations, thus leading to improved performance, as shown in Fig. 2. Training and inference strategies are presented in §3.6.
@@ 3.1 Preliminaries
SAM2 [39] is a pioneering vision foundation model designed for promptable visual segmentation tasks. Different from SAM [19], SAM2 includes a memory module that stores information about the object from previous frames. It contains an image encoder, memory attention, prompt encoder, mask decoder, memory encoder, and memory bank. For each frame, the image encoder extracts representative visual features, which are then conditioned on the features and predictions of past frames. If a prompt (point, box, or mask) is given, the prompt encoder encodes it into sparse or dense embeddings. Exploiting memory-conditioned features and prompt embeddings, the mask decoder outputs the segmentation mask. The memory encoder then updates the memory bank with the output mask and the unconditioned frame embedding to support the segmentation of subsequent frames. SAM2 is pre-trained on SA-1B [19] and further trained on SA-V [39], achieving strong performance across video and image segmentation tasks. For more details, please refer to [39].
@@ 3.2 Decamouflaged Token
Given a video clip containing $m$ frames, we denote all frames as $\{\mathbf{I}_{t-m+1},\cdots,\mathbf{I}_{i},\cdots,\mathbf{I}_{t}\}$ with ground-truth segmentation masks of $\{\mathbf{S}_{t-m+1},\cdots,\mathbf{S}_{i},\cdots,\mathbf{S}_{t}\}$. Especially, $\mathbf{I}_{t}$ is the current frame for the purpose of easy explanation. We use the image encoder to extract features for all frames, denoted as $\{\mathbf{F}_{t-m+1},\cdots,\mathbf{F}_{i},\cdots,\mathbf{F}_{t}\}$. Here, $\mathbf{F}_{i}$ can be further represented as $\{\mathbf{F}_{i}^{0},\cdots,\mathbf{F}_{i}^{j},\cdots,\mathbf{F}_{i}^{L-1}\}$, containing feature maps extracted from $L$ different intermediate layers or blocks, where $\mathbf{F}_{i}^{j}\in\mathbb{R}^{c_{j}\times h_{j}\times w_{j}}$, with $c_{j}$, $h_{j}$, and $w_{j}$ representing channels, height, and width, respectively.
SAM2’s output tokens include an object score (occlusion) token, an IoU token, and mask tokens. To enhance SAM2’s ability of segmenting camouflaged objects, we introduce a new learnable decamouflaged token $\mathbf{T}\in\mathbb{R}^{1\times 256}$, enabling it to optimize features for segmenting camouflaged objects. As depicted in Fig. 2, integrated with SAM2’s output tokens, the decamouflaged token undergoes the same layers as output tokens within SAM2’s mask decoder: two attention blocks including self-attention, followed by token-to-image (T2I) and image-to-token (I2T) cross attention, and the last T2I cross attention. After this, the output decamouflaged token is denoted as $\mathbf{T}^{\prime}$. This token is updated through back-propagated gradients, while SAM2’s weights remain frozen. $\mathbf{T}^{\prime}$ is then passed through an MLP layer to participate in computing CamSAM2’s final mask logits, which will be explained in §3.4.

### v3

@@ Abstract
Video camouflaged object segmentation (VCOS), aiming at segmenting camouflaged objects that seamlessly blend into their environment, is a fundamental vision task with various real-world applications. With the release of SAM2, video segmentation has witnessed significant progress. However, SAM2’s capability of segmenting camouflaged videos is suboptimal, especially when given simple prompts such as point and box. To address the problem, we propose Camouflaged SAM2 (CamSAM2), which enhances SAM2’s ability to handle camouflaged scenes without modifying SAM2’s parameters. Specifically, we introduce a decamouflaged token to provide the flexibility of feature adjustment for VCOS. To make full use of fine-grained and high-resolution features from the current frame and previous frames, we propose implicit object-aware fusion (IOF) and explicit object-aware fusion (EOF) modules, respectively. Object prototype generation (OPG) is introduced to abstract and memorize object prototypes with informative details using high-quality features from previous frames. Extensive experiments are conducted to validate the effectiveness of our approach. While CamSAM2 only adds negligible learnable parameters to SAM2, it substantially outperforms SAM2 on three VCOS datasets, especially achieving 12.2 mDice gains with click prompt on MoCA-Mask and 19.6 mDice gains with mask prompt on SUN-SEG-Hard, with Hiera-T as the backbone. The code is available at https://github.com/zhoustan/CamSAM2.
@@ 1 Introduction
Camouflaged object detection (COD) and video camouflaged object segmentation (VCOS) aim to identify objects that blend seamlessly into their surroundings. Unlike standard object segmentation tasks, where objects typically exhibit clear boundaries and contrast with the background, camouflaged objects are naturally indistinguishable from the background. These tasks have various applications in wildlife monitoring, surveillance, and search-and-rescue operations [roy2022computer, stephenson2020technological]. COD focuses on detecting camouflaged objects in individual images, while VCOS extends it to video sequences, adding the complexity of modeling temporal information across frames. Despite recent advancements in COD [pang2024zoomnext, cheng2022implicit, zhang2024learning, chen2024sam, pang2024open, zhao2024focusdiffuser, yin2024camoformer] and VCOS [pang2024zoomnext, cheng2022implicit, hui2024endow, meeran2024sam, Kowal_2022_CVPR, bideau2024right, lu2024weakly], the performance remains far from satisfactory compared to general segmentation tasks.
@@
The recently introduced vision foundation model, Segment Anything Model 2 (SAM2) [ravi2025sam], marks a significant advancement in video segmentation. SAM2 has learned rich and generalizable representations for natural scenes on the SA-1B [kirillov2023segment] (11M images, 1B masks) and SA-V [ravi2025sam] (50.9K videos, 35.5M masks) datasets. Therefore, its features are optimized for natural scenes, while SAM2’s ability of segmenting camouflaged objects is suboptimal, as in [zhou2025sam2, tang2024evaluatingsam2srolecamouflaged]. Following the setting of SAM2, we apply point, box, or mask prompts on the first frame of each video. Specifically, we randomly sample a point within the ground truth region as a point prompt, extract the tight bounding box as the box prompt, or use the full ground-truth mask as the mask prompt. As shown in Fig.˜1, with a point prompt, SAM2 segments only part of a camouflaged animal (hedgehog), indicating that there is still room for improvements in VCOS.
Figure 1: Illustration of SAM2 and CamSAM2. Top: SAM2’s segmentation of the camouflaged object is suboptimal, primarily because its feature optimization is biased toward natural videos, and its design does not account for the unique challenges inherent to VCOS. Bottom: CamSAM2 improves SAM2’s ability to segment and track camouflaged objects by introducing a decamouflaged token, IOF to enhance features with high-resolution features, and EOF and OPG to further enhance features by exploiting informative object details across time. CamSAM2 only adds a limited number of parameters to SAM2 while keeping all SAM2’s parameters fixed and fully inheriting SAM2’s zero-shot ability. The segmentation result is overlaid in orange on the frame.
@@
This paper aims to develop a model for accurate segmentation in camouflaged videos, requiring both natural image understanding and effective identification of camouflaged objects in complex environments. To achieve this, we identify two core challenges in adapting SAM2 for VCOS: (1) SAM2 is optimized for natural scenes rather than camouflaged environments. (2) The architecture does not account for the complexities of segmenting and tracking camouflaged objects across time. For VCOS, accurately segmenting camouflaged objects for a frame requires: a) exploiting fine-grained and detailed features from the frame, and b) considering the temporal evolvement of fine-grained features from previous frames. For exploiting temporal information, SAM2 is equipped with a memory module containing a memory encoder and a memory bank. However, only low-resolution and coarse features are encoded into the memory, which is suboptimal for accurate VCOS.
@@
To tackle the above limitations and fully keep SAM2’s ability to process natural videos, we introduce Camouflaged SAM2, dubbed as CamSAM2, equipping SAM2 with the ability to effectively tackle VCOS, as depicted in Fig.˜1. CamSAM2 includes a learnable decamouflaged token, which extends the token structure of SAM2 and provides flexibility to optimize features for VCOS without modifying SAM2’s trained parameters. To exploit the fine-grained features of the frame, we propose the Implicit Object-aware Fusion (IOF) module, which leverages high-resolution features from the early layers of the image encoder to enhance the model’s perception of fine-grained details. To make use of detailed features from previous frames, we further propose Object Prototype Generation (OPG) to abstract high-quality features within the object region into informative object prototypes through Farthest Point Sampling (FPS) and k-means. Those object prototypes are saved to memory for easy usage by the Explicit Object-aware Fusion (EOF) module that is designed to integrate explicit object-aware information across the temporal dimension. Our design avoids saving the high-resolution features in the memory and only adds negligible computations to SAM2 while accounting for a large amount of temporal information.
@@
We conduct extensive experiments in §4 to validate the effectiveness of CamSAM2 on three VCOS benchmarks: two camouflaged animal datasets, MoCA-Mask [cheng2022implicit] and CAD [bideau2016s], and one camouflaged medical dataset, SUN-SEG [ji2022video]. Our experiments show that CamSAM2 significantly outperforms SAM2 by achieving improvements of 12.2/13.1 mDice scores with click prompt on MoCA-Mask for Hiera-T/Hiera-S backbones, and 19.6 mDice gains with mask prompt on SUN-SEG-Hard for Hiera-T backbone. When directly evaluating CamSAM2 on CAD without further finetuning, we observe strong zero-shot ability. Since all SAM2’s weights remain unchanged, CamSAM2 totally inherits SAM2’s capability on segmenting natural videos. In summary, our contributions are three-fold:
@@
(1) We propose CamSAM2 to equip SAM2 with the ability to segment and track camouflaged objects in videos while keeping SAM2’s strong generalizability in natural videos. (2) CamSAM2 introduces a decamouflaged token to achieve easy feature adjustments for the VCOS task without affecting SAM2’s trained weights. To effectively exploit the crucial fine-grained and high-resolution features from both the current frame and previous frames, we propose IOF, EOF, and OPG modules. (3) Our approach clearly outperforms SAM2 and sets new state-of-the-art performance on public VCOS datasets. Experiments also show the strong zero-shot ability of CamSAM2 in the domain of VCOS.
@@ 2.1 Camouflaged object detection
Camouflaged Scene Understanding (CSU) focuses on interpreting scenes where objects blend closely with their backgrounds, such as natural environments like forests, oceans, and deserts. Early works in this field primarily involved the collection of extensive image and video datasets, such as CAMO [le2019anabranch], COD10K [fan2020camouflaged], NC4K [yunqiu_cod21], CAD [bideau2016s], MoCA-Mask [cheng2022implicit], and MoCA-Mask-Pseudo [cheng2022implicit], which laid the foundation for CSU. Traditional COD methods extract foreground-background features using optical features [beiderman2010optical], color, and texture [kavitha2011efficient]. Deep learning has advanced COD with CNNs and transformers. SINet [fan2020camouflaged] and SINet-V2 [fan2021concealed] enhance fine-grained cues by applying receptive fields and texture-enhanced modules, while DQNet [sun2022dqnet] applies cross-modal detail querying to detect subtle features. Transformer-based models like CamoFormer [yin2024camoformer] leverage multi-scale feature extraction with masked separable attention, and WSSCOD [zhang2024learning] employs a frequency transformer and noisy pseudo labels for weak supervision. ZoomNeXt [pang2024zoomnext] further optimizes multi-scale extraction via a collaborative pyramid network. These advancements refine COD by integrating sophisticated architectures and diverse supervision strategies.
@@ 2.2 Video camouflaged object segmentation
VCOS [lu2024weakly, yu2024tokenmotion, Kowal_2022_CVPR, xie2022segmenting] extends COD to videos, introducing challenges from motion, dynamic backgrounds, and temporal consistency. Former VCOS models tackle these with motion learning, spatial-temporal attention, and advanced segmentation techniques to maintain object coherence across frames. Motion-guided models enhance segmentation by leveraging motion cues. IMEX [hui2024implicit] integrates implicit and explicit motion learning for robust detection. TMNet [yu2024tokenmotion] refines motion features with a transformer-based encoder and neighbor connection decoder. Flow-SAM [xie2024flowsam] uses optical flow as input or a prompt, guiding SAM to detect moving camouflaged objects. Spatial-temporal attention enhances the tracking of camouflaged objects. TSP-SAM [hui2024endow] and SAM-PM [meeran2024sam] improve SAM’s ability to detect subtle movements. Static-Dynamic-Interpretability [Kowal_2022_CVPR] quantifies static and dynamic information in spatial-temporal models, aiding balanced approaches. Assessing camouflage quality is also essential for VCOS. CAMEVAL [lamdouar2023making] introduces scores evaluating background similarity and boundary visibility, refining datasets and improving model robustness. These advancements drive more accurate and effective VCOS systems.
@@ 2.3 Segment Anything Model 2
SAM2 [ravi2025sam] is a vision foundation model for promptable segmentation across images and videos. Compared to SAM [kirillov2023segment], which is limited to image segmentation, SAM2 offers a significant performance leap in video segmentation. SAM2 has demonstrated strong capabilities in many tasks, including medical image, video and 3D segmentation [mansoori2025polyp, shen2025performance, yu2024sam2roboticsurgery, liu2024surgical, he2024shortreviewevaluationsam2s, chen2024sam2adapterevaluatingadapting, xiong2024sam2unetsegment2makes, yu2024noveladaptationvideosegmentation], video object tracking and segmentation [stanczyk2024masks, pan2024videoobjectsegmentationsam], remote sensing [rafaeli2024promptbasedsegmentationmultipleresolutions], 3D mesh and point cloud segmentation [tang2024segmentmeshzeroshotmesh], COD and VCOS [zhou2025sam2, chen2024sam2adapterevaluatingadapting, xiong2024sam2unetsegment2makes, tang2024evaluatingsam2srolecamouflaged]. In previous works [zhou2025sam2, tang2024evaluatingsam2srolecamouflaged], although SAM2’s performance on camouflaged video segmentation has surpassed most existing methods, these studies primarily focused on direct evaluation, lightweight fine-tuning, or integrating SAM2 with multimodal large language models. However, they did not address the fundamental architectural challenges of adapting SAM2 to VCOS, leaving a significant performance gap compared to other VOS tasks, especially when using simple prompts.
@@ 3 Method
Figure 2: Overall architecture of CamSAM2. CamSAM2 effectively captures and segments camouflaged objects by leveraging implicit and explicit object-aware information from the current or previous frames. It includes the following key components: (a) the decamouflaged token, which extends SAM2’s token structure to learn features suitable for camouflaged objects; (b) an IOF module to enrich memory-conditioned features with implicitly object-aware high-resolution features; (c) an EOF module to aggregate explicit object-aware features; and (d) an OPG module, generating informative object prototypes, which guides cross-attention in EOF. These components work together to preserve fine details, enhance segmentation quality, and track camouflaged objects across time.
We propose CamSAM2, equipping SAM2 with the ability to accurately segment camouflaged objects in videos while retaining SAM2’s original capabilities. §3.1 briefly reviews the architecture of SAM2. From §3.2 to §3.5, we describe CamSAM2 tailored for VCOS. With fixing SAM2’s parameters, CamSAM2 proposes a learnable decamouflaged token, Implicit and Explicit Object-aware Fusion, and Object Prototype Generation to enhance feature representations, thus leading to improved performance, as shown in Fig.˜2.
@@ 3.1 Preliminaries
SAM2 [ravi2025sam] is a pioneering vision foundation model designed for promptable visual segmentation tasks. Different from SAM [kirillov2023segment], SAM2 includes a memory module that stores information about the object from previous frames. It contains an image encoder, memory attention, prompt encoder, mask decoder, memory encoder, and memory bank. For each frame, the image encoder extracts representative visual features, which are then conditioned on the features and predictions of past frames. If a point or box prompt is given, the prompt encoder encodes it into sparse or dense embeddings, then segments the prompted frame; if a mask prompt is given, SAM2 directly uses the mask as the current frame output. Exploiting memory-conditioned features and prompt embeddings, the mask decoder outputs the segmentation mask. The memory encoder then updates the memory bank with the output mask and the unconditioned frame embedding to support the segmentation of subsequent frames. SAM2 is pre-trained on SA-1B [kirillov2023segment] and further trained on SA-V [ravi2025sam], achieving strong performance across video and image segmentation tasks. For more details, please refer to [ravi2025sam].
@@ 3.2 Decamouflaged token
Given a video clip containing $m$ frames, we denote all frames as $\{\mathbf{I}_{t-m+1},\cdots,\mathbf{I}_{i},\cdots,\mathbf{I}_{t}\}$ with ground-truth segmentation masks of $\{\mathbf{S}_{t-m+1},\cdots,\mathbf{S}_{i},\cdots,\mathbf{S}_{t}\}$. Especially, $\mathbf{I}_{t}$ is the current frame for the purpose of easy explanation. We use the image encoder to extract features for all frames, denoted as $\{\mathbf{F}_{t-m+1},\cdots,\mathbf{F}_{i},\cdots,\mathbf{F}_{t}\}$. Here, $\mathbf{F}_{i}$ can be further represented as $\{\mathbf{F}_{i}^{0},\cdots,\mathbf{F}_{i}^{j},\cdots,\mathbf{F}_{i}^{L-1}\}$, containing feature maps extracted from $L$ different intermediate layers, where $\mathbf{F}_{i}^{j}\in\mathbb{R}^{c_{j}\times h_{j}\times w_{j}}$, with $c_{j}$, $h_{j}$, and $w_{j}$ representing channels, height, and width, respectively.
SAM2’s output tokens include an object score (occlusion) token, an IoU token, and mask tokens. To enhance SAM2’s ability of segmenting camouflaged objects, we introduce a new learnable decamouflaged token $\mathbf{T}\in\mathbb{R}^{1\times 256}$, enabling it to optimize features for segmenting camouflaged objects. As depicted in Fig.˜2, integrated with SAM2’s output tokens, the decamouflaged token undergoes the same layers as output tokens within SAM2’s mask decoder.
After this, the output decamouflaged token is denoted as $\mathbf{T}^{\prime}$. This token is updated through back-propagated gradients, while SAM2’s weights remain frozen. $\mathbf{T}^{\prime}$ is then passed through an MLP layer to participate in computing CamSAM2’s final mask logits, which will be explained in §3.4.